# Entropy Meets Importance: A Unified Head Importance–Entropy Score for Stable and Efficient Transformer Pruning

### v1

@@ D.2 Heatmap of Importance Scores and Pruning Results
D.3 3D Analysis of Attention Head Importance Scores
@@
D.3.1 Difference in Pruning Patterns
Pruning heatmaps (Fig. 9) reveal systematic differences between the methods. HIS-based pruning tends to remove heads primarily from the lower layers, producing an approximately bottom-up pattern consistent with its one-step gradient saliency. In contrast, HIES yields a more dispersed selection spanning lower, middle, and upper layers. We attribute this to the entropy-aware term, which leverages structural properties of the attention distribution (concentration vs. dispersion) in addition to gradient sensitivity, thereby promoting diversity across layers in pruning decisions.
D.3.2 Performance Inversion Across Sparsity Regimes
@@
In the early pruning phase, the model contains a substantial number of redundant heads. Here, gradient-based importance scores (HIS) are sufficient to identify and remove low-sensitivity heads, as they reflect the immediate (one-step) loss change. Consequently, HIS performs slightly better than HIES in both accuracy and stability under light pruning.
@@
D.4 Experimental Results on Downstream Tasks
Fig. 10 reports downstream evaluations of HIES versus the HIS baseline on CIFAR-100, Food-101, and Fashion-MNIST. Across all three benchmarks, HIES consistently sustains higher accuracy under aggressive pruning, whereas HIS exhibits rapid degradation once the pruning ratio exceeds 20%. On CIFAR-100, HIS collapses beyond moderate sparsity, while HIES exhibits slower degradation and retains substantially higher accuracy relative to HIS even at 40–50%. Food-101 reveals a similar trend, with HIES delivering substantial and consistent gains over HIS across all pruning levels. On Fashion-MNIST, HIS undergoes steep drops after 20% pruning, in contrast to the stable performance of HIES up to 50%. These results demonstrate that HIES reliably mitigates sharp-drop phenomena and delivers robust, stable improvements over HIS across heterogeneous downstream tasks.
@@ D.4 Experimental Results on Downstream Tasks
D.5 Sensitivity Analysis - Ablation on $\alpha$
@@
We sweep the mixing coefficient $\alpha\in[0,1)$ that interpolates the gradient-based head-importance (HIS) and attention-entropy (AE) signals in HIES,
@@
As expected, larger $\alpha$ upweights HIS and preserves heads with strong task relevance, whereas smaller $\alpha$ upweights AE and retains low-entropy, focused heads. We choose a single $\alpha^{\star}$ on a held-out validation split and fix it for all reported experiments; the resulting accuracy–sparsity profiles are shown in Fig. 11.

### v2

@@ D.4 Heatmap of Importance Scores and Pruning Results
D.5 3D Analysis of Attention Head Importance Scores
@@
D.5.1 Difference in Pruning Patterns
As shown in Figure 9, Pruning heatmaps reveal systematic differences between the methods. HIS-based pruning tends to remove heads primarily from the lower layers, producing an approximately bottom-up pattern consistent with its one-step gradient saliency. In contrast, HIES yields a more dispersed selection spanning lower, middle, and upper layers. We attribute this to the entropy-aware term, which leverages structural properties of the attention distribution (concentration vs. dispersion) in addition to gradient sensitivity, thereby promoting diversity across layers in pruning decisions.
D.5.2 Performance Inversion Across Sparsity Regimes
@@
In the early pruning phase, the model contains a substantial number of redundant heads. Here, gradient-based HIS is sufficient to identify and remove low-sensitivity heads, as they reflect the immediate (one-step) loss change. Consequently, HIS performs slightly better than HIES in both accuracy and stability under light pruning.
@@
D.6 Experimental Results on Large-scale Reasoning Tasks
Figure 10 reports reasoning performance on GSM8K and MMLU with $\text{LLaMA-2}_{\text{7B}}$, comparing HIES against HIS-based pruning. Across both benchmarks, HIES consistently outperforms HIS across pruning regimes, achieving an average accuracy improvement of 14.67% and demonstrating improved robustness on reasoning tasks.
Figure 10: Results on GSM8K (math word-problem reasoning) and MMLU (10-task knowledge reasoning) with $\text{LLaMA-2}_{\text{7B}}$.
D.7 Experimental Results on Downstream Tasks
Figure 11 reports downstream evaluations of HIES versus the HIS baseline on CIFAR-100, Food-101, and Fashion-MNIST. Across all three benchmarks, HIES consistently sustains higher accuracy under aggressive pruning, whereas HIS exhibits rapid degradation once the pruning ratio exceeds 20%. On CIFAR-100, HIS collapses beyond moderate sparsity, while HIES exhibits slower degradation and retains substantially higher accuracy relative to HIS even at 40–50%. Food-101 reveals a similar trend, with HIES delivering substantial and consistent gains over HIS across all pruning levels. On Fashion-MNIST, HIS undergoes steep drops after 20% pruning, in contrast to the stable performance of HIES up to 50%. These results demonstrate that HIES reliably mitigates sharp-drop phenomena and delivers robust, stable improvements over HIS across heterogeneous downstream tasks.
@@ D.7 Experimental Results on Downstream Tasks
D.8 Sensitivity Analysis - Ablation on $\alpha$
@@
We sweep the mixing coefficient $\alpha\in[0,1)$ that interpolates the gradient-based HIS and AE signals in HIES,
@@
As expected, larger $\alpha$ upweights HIS and preserves heads with strong task relevance, whereas smaller $\alpha$ upweights AE and retains low-entropy, focused heads. We choose a single $\alpha^{\star}$ on a held-out validation split and fix it for all reported experiments; the resulting accuracy–sparsity profiles are shown in Figure 12.
D.9 Combining Attention Entropy with Other Importance Signals
To examine the applicability of AE beyond gradient-based scores, we combine it with a different importance signal based on the L2 norm of attention outputs. This experiment assesses the generality of AE and verifies that it can serve as a complementary stabilization term under different importance signals.
Figure 13 reports the impact of combining AE with the L2-norm–based importance score on $\text{LLaMA-2}_{\text{7B}}$ across five benchmarks: HellaSwag, Winogrande, ARC-e, ARC-c, and OBQA. Across all benchmarks, incorporating AE consistently mitigates performance degradation under aggressive pruning, indicating that the stabilizing effect of AE is not restricted to a particular importance formulation.
Figure 13: Accuracy and stability improvements when combining AE with gradient L2-norm–based importance scores, compared to using the L2-norm alone. We conduct experiments on $\text{LLaMA-2}_{\text{7B}}$ for HellaSwag, Winogrande, PIQA, ARC-e, ARC-c, and OBQA.
Appendix E Efficiency Analysis
E.1 Computational Efficiency and FLOPs Reduction
As the pruning ratio increases, the total FLOPs decrease approximately linearly: removing 10% of attention heads yields an $\approx$4% reduction in FLOPs, while pruning 50% of heads achieves an $\approx$20% reduction. Under HIS-only pruning, model accuracy on TinyBERT with SST-2 sharply degrades beyond a 42% pruning ratio, at which point only an $\approx$16% FLOPs reduction can be attained without critical performance loss. In contrast, HIES-based pruning maintains at least 80% validation accuracy up to a 60% pruning ratio, corresponding to an $\approx$23% FLOPs reduction relative to the original model. Extending the feasible pruning regime from 42% to 60% therefore delivers an additional $\approx$7 percentage-point reduction in FLOPs, demonstrating that HIES enables substantially greater computational efficiency without compromising task performance.
E.2 Runtime Overhead and Implementation Details
The above efficiency gains are achieved with only a small computational overhead. HIES is computed using lightweight forward hooks attached to the attention modules. Specifically, HIS reuses gradients produced during standard backpropagation, while AE is obtained directly from attention probabilities in the forward pass. Both signals are collected via non-intrusive forward hooks without modifying model internals or introducing additional forward or backward passes. As a result, the runtime overhead is minimal—typically on the order of a few seconds— with negligible additional memory cost beyond a small buffer for storing per-head statistics. Despite this low overhead, HIES consistently outperforms baselines that rely on full fine-tuning, achieving superior efficiency in both computation and accuracy.